[figure: 9.7 cm]

Electronic structure of Ce_{2}TSi_{3} (T=Rh,Pd) compounds

A. Szytuła^{a}, A. Jezierski^{b} and B. Penc^{a} aInstutute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków bInstitute of Molecular Physics, Polish Academy of Sciences, 60-179 Poznań

Electronic structure of the ternary Ce_{2}TSi_{3} (T=Rh, Pd) compounds wich crystalize in the hexagonal AlB_{2} structure was studied by X-ray photoelectron spectroscopy (XPS). Core-levels and valence bands were investigated. The XPS valence bands are compared with the ones calculated using TB LMTO method. The photoemission measurements were performed using a commercial LHS 10 SPECS spectrometer with the following radiation sources: nonmonochromatized MgK_{α} (hν=1253.6 eV), monochromatized and non-monochromatized AlK_{α} (hν=1486.6 eV). The measurements were made at room temperature in high vacuum (2·10^{−9} mbar). The VB XPS spectra of Ce_{2}RhSi_{3} and Ce_{2}PdSi_{3} are established mainly by 4d states of Rh and Pd- atoms. The XPS spectra of 3d Ce consists of three peaks that correspond to the f^{0}, f^{1} and f^{2} configurations. The gap between energies of the f^{1} states corresponding to the energy levels of 3d_{5/2} and 3d_{3/2} states give a spin-orbit splitting Δ_{SO}, which is equal to 18.7 eV for Ce_{2}PdSi_{3} and 18.5 eV for Ce_{2}RhSi_{3}. Based on the Gunnarsson-Schönhammer theoretical model, the intensity ratio f^{0}/(f^{0} + f^{1} + f^{2}) gives the f-occupation probability in the final state. The obtained results suggest that about 0.25 ions have the f^{0} configuration. The shift of the shake down satellites in relation to the main peak is about 3.3 eV in Ce_{2}PdSi_{3} and 3.6 eV in Ce_{2}RhSi_{3}. The intensity ratios r = f^{2}/(f^{1} + f^{2}) for Ce_{2}PdSi_{3} and Ce_{2}RhSi_{3} are near 0.5. Such r values correspond to a very large value of Δ(about 300 meV). Similar, large Δ’s (about 220 eV) are observed in Ce stannides and antimonides of the 1:2:2 type.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Szytuła

Address for correspondence : Instutute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków

Email address : szytula@if.uj.edu.pl